# Quantum Dynamic Programming

### v1

@@ IV.2.1 Distance bounds after a DME implementation
Let $\rho=(1-x)\psi+x\frac{\mathmybb{1}_{\rm rel}}{2}$ be a state with an associated pure state $\psi$ and a mixedness parameter $x\in[0,\varepsilon]$ for some $10\leq\varepsilon<$ 1. A dynamic unitary channel
$\hat{\mathbf{\Gamma}}^{(\rho)}_{L}=\hat{\mathbf{E}}^{(\rho)}_{\alpha_{L}}\circ\hat{\mathbf{E}}^{(\tau)}_{\beta_{L}}\circ\cdots\circ\hat{\mathbf{E}}^{(\rho)}_{\alpha_{1}}\circ\hat{\mathbf{E}}^{(\tau)}_{\beta_{1}}$ (77)
that uses the same angles $\{\alpha_{l},\beta_{l}\}_{l}$ used in $\hat{\mathbf{\Gamma}}^{(\psi)}_{L}$ has the distance
$\frac{1}{2}\left\lVert\hat{\mathbf{\Gamma}}^{(\psi)}_{L}-\hat{\mathbf{\Gamma}}^{(\rho)}_{L}\right\rVert_{\rm Tr}\leq\frac{L\pi\varepsilon}{2}\ ,$ (78)
@@
$\left\lVert\hat{\mathbf{\Gamma}}^{(\psi)}_{L}-\hat{\mathbf{\Gamma}}^{(\rho)}_{L}\right\rVert_{\rm Tr}\leq\sum_{i=1}^{L}\left\lVert\hat{\mathbf{E}}^{(\psi)}_{\alpha_{i}}-\hat{\mathbf{E}}^{(\rho)}_{\alpha_{i}}\right\rVert_{\rm Tr}\ ,$ (79)
from the triangle inequality and the unitary invariance of the trace norm. The distance between $\hat{\mathbf{E}}^{(\rho)}_{\alpha}$ and $\hat{\mathbf{E}}^{(\psi)}_{\alpha}$ can be written as
$\frac{1}{2}\left\lVert\hat{\mathbf{E}}^{(\psi)}_{\alpha}-\hat{\mathbf{E}}^{(\rho)}_{\alpha}\right\rVert_{\rm Tr}=\frac{1}{2}\max_{\sigma\in\mathcal{S}(\mathcal{H}_{\rm rel})}\left\lVert\hat{\mathbf{E}}^{(\psi)}_{\alpha}\left(\sigma\right)-\hat{\mathbf{E}}^{(\rho)}_{\alpha}(\sigma)\right\rVert_{1}=\frac{1}{2}\max_{\sigma\in\mathcal{S}(\mathcal{H}_{\rm rel})}\operatorname{Tr}\left\{\sqrt{\left(\hat{\mathbf{E}}^{(\psi)}_{\alpha}\left(\sigma\right)-\hat{\mathbf{E}}^{(\rho)}_{\alpha}(\sigma)\right)^{2}}\right\}\ .$ (80)
@@
$\psi=\begin{bmatrix}1&0\\ 0&0\end{bmatrix},\quad\rho=\begin{bmatrix}1-\frac{x}{2}&0\\ 0&\frac{x}{2}\end{bmatrix},\quad\sigma=\begin{bmatrix}p&q\\ q^{*}&1-p\end{bmatrix}\ ,$ (87)
@@
$\left(\hat{\mathbf{E}}^{(\psi)}_{\alpha}(\sigma)-\hat{\mathbf{E}}^{(\rho)}_{\alpha}(\sigma)\right)^{2}$ $=\begin{bmatrix}4\lvert q\rvert^{2}\sin^{2}\left(\frac{\alpha x}{2}\right)&0\\ 0&4\lvert q\rvert^{2}\sin^{2}\left(\frac{\alpha x}{2}\right)\end{bmatrix}\ .$ (90)
The state $\sigma$ giving the maximum distance is the one having the maximal $\lvert q\rvert$. The condition that $\sigma\geq 0$ imposes $\lvert q\rvert\leq\frac{1}{2}$, and $\frac{1}{2}\lVert\hat{\mathbf{E}}^{(\psi)}_{\alpha}-\hat{\mathbf{E}}^{(\rho)}_{\alpha}\rVert_{\rm Tr}=\lvert\sin\left(\frac{\alpha x}{2}\right)\rvert$. From $\alpha\in[-\pi,\pi]$ and $\lvert\alpha\varepsilon\rvert<\pi$, we bound $\lvert\sin(\frac{\alpha x}{2})\rvert<\frac{\pi\varepsilon}{2}$ and achieve Eq. (78). ∎
@@
$\delta_{n}\coloneq\left(T_{(2L+1)}\left(\delta_{n-1}^{-1}\right)\right)^{-1}+\epsilon\ ,$ (91)
and use it as the parameter $q=\delta_{n}$ when choosing angles Eq. (58). At the same time, we also demonstrate that $\delta_{n}$ bounds the distance to the target state, after an iteration of DME implementation with error $\epsilon$, i.e. it is the counterpart to Eq. (60) used for the exact iteration. Precise statements of this result are delineated in Lemma 14 and Corollary 15. Further analysis on Eq. (91) itself will be made in Sec. IV.3.
@@
We consider iteration $n$ of the nested Grover search algorithm. Assume that the initial state for the iteration $\rho_{n-1}=(1-x_{n-1})\psi_{n-1}+x_{n-1}\frac{\mathmybb{1}_{\rm rel}}{2}$ is unknown to us but satisfies the following:
@@
  • The mixedness parameter is bounded by $x_{n-1}\leq\frac{\epsilon}{L\pi}$ for some number $\epsilon$.
Then, it is possible to prepare a state $\rho^{\prime}_{n}=(1-x_{n}^{\prime})\psi_{n}+x^{\prime}\frac{\mathmybb{1}_{\rm rel}}{2}$, such that
$\frac{1}{2}\lVert\tau-\rho^{\prime}_{n}\rVert_{1}$ $\leq\delta_{n}-\frac{\epsilon}{4},$ (92)
where $\delta_{n}$ is defined in Eq. (91). The preparation requires consuming $\mathcal{O}(L^{2}\epsilon^{-1})$ copies of $\rho_{n-1}$.
@@
The preparation is performed by implementing nested fixed-point Grover search with DME memory-usage queries. Firstly, we choose the angles $\{\alpha_{l}(q),\beta_{l}(q)\}_{l}$ for the dynamic unitary $\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}$. Following Lemma 10, we set $q=\operatorname{sech}((2L+1)\operatorname{arcsech}(\delta_{n-1}))$, or equivalently, $q=\delta_{n}-\epsilon$. Secondly, we denote the total number of DME queries ^s$\hat{\mathcal{E}}^{(\rho_{n-1})}_{s}$ used for implementing $\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}$ as $M$. Since $L$ memory-calls (partial reflections around $\rho_{n-1}$) are contained in $\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}$, each memory-call employs $\frac{M}{L}$ DME memory-usage queries. The implemented channel reads
$\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}\coloneq\left(\hat{\mathcal{E}}^{(\rho_{n-1})}_{s_{L}}\right)^{\frac{M}{L}}\circ\hat{\mathbf{E}}^{(\tau)}_{\beta_{L}}\circ\cdots\circ\left(\hat{\mathcal{E}}^{(\rho_{n-1})}_{s_{1}}\right)^{\frac{M}{L}}\circ\hat{\mathbf{E}}^{(\tau)}_{\beta_{1}}\ ,$ (93)
@@
$\rho^{\prime}_{n}:=\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}\left(\rho_{n-1}\right)=(1-x_{n}^{\prime})\psi_{n}+x_{n}^{\prime}\frac{\mathmybb{1}_{\rm rel}}{2}\ .$ (94)
We additionally define the state after an exact unitary channel $\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}$ as
$\tilde{\rho}_{n}:=\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}\left(\rho_{n-1}\right)=(1-x_{n-1})\phi_{n}+x_{n-1}\frac{\mathmybb{1}_{\rm rel}}{2}\ .$ (95)
Since $\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}$ is a unitary channel, it leaves the projector $\mathmybb{1}_{\rm rel}$ intact in Eq. (95), and the associated pure state
$\phi_{n}=\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})\ ,$ (96)
@@
    $\lVert\tau-\phi_{n}\rVert_{1}=\left\lVert\tau-\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}\leq\left\lVert\tau-\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}+\left\lVert\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}(\psi_{n-1})-\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}\ .$ (97)
    The second term in the RHS is bounded by the trace norm distance of channels $\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}$ and $\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}$. Applying Proposition 13 and using the assumption $x_{n-1}\leq\frac{\epsilon}{L\pi}$,
    $\left\lVert\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}(\psi_{n-1})-\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}\leq\left\lVert\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}-\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}\right\rVert_{\rm Tr}\leq\epsilon\ .$ (98)
    The other term $\lVert\tau-\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}(\psi_{n-1})\rVert_{1}$ resembles the final distance after an exact nested fixed-point Grover search in Lemma 10. However, we cannot directly apply Lemma 10, since we do not know the initial distance $\lVert\tau-\psi_{n-1}\rVert_{1}$ exactly; we only assume that it is bounded by $\delta_{n-1}$. Hence, we must start from the more general relation Eq. (59), which gives the equality
    $\frac{1}{2}\left\lVert\tau-\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}=(\delta_{n}-\epsilon)T_{2L+1}\left(T_{1/(2L+1)}\left((\delta_{n}-\epsilon)^{-1}\right)\frac{1}{2}\left\lVert\tau-\psi_{n-1}\right\rVert_{1}\right)\ ,$ (99)
    substituting $q=\delta_{n}-\epsilon$. Recalling that $\delta_{n}-\epsilon=(T_{(2L+1)}(\delta_{n-1}^{-1}))^{-1}$ and Eq. (55), we get $T_{1/(2L+1)}((\delta_{n}-\epsilon)^{-1})=\delta_{n-1}^{-1}$. Eq. (99) then becomes
    $\frac{1}{2}\left\lVert\tau-\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}=(\delta_{n}-\epsilon)T_{2L+1}\left(\delta_{n-1}^{-1}\frac{1}{2}\left\lVert\tau-\psi_{n-1}\right\rVert_{1}\right)\ .$ (100)
    Chebyshev polynomials follow $T_{2L+1}(x)\leq 1$ whenever $x\in[0,1]$, and the argument of $T_{2L+1}$ in the RHS, $\delta_{n-1}^{-1}\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}$ is indeed smaller than or equal to $1$ from the assumption of the Lemma. Then $\frac{1}{2}\lVert\tau-\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}(\psi_{n-1})\rVert_{1}\leq\delta_{n}-\epsilon$ follows, and we finally obtain the bound
    $\frac{1}{2}\lVert\tau-\phi_{n}\rVert_{1}\leq\delta_{n}-\frac{\epsilon}{2}\ .$ (101)
@@
    $\lVert\phi_{n}-\rho^{\prime}_{n}\rVert_{1}$ $\leq\left\lVert\phi_{n}-\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}+\left\lVert\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})-\rho^{\prime}_{n}\right\rVert_{1}$
    $=\left\lVert\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}+\left\lVert\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\rho_{n-1})\right\rVert_{1}\ .$ (102)
@@
    $\left\lVert\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}\leq\left\lVert\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}-\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}\right\rVert_{\rm Tr}\leq\sum_{l=1}^{L}\frac{M}{L}\left\lVert\hat{\mathbf{E}}^{(\rho_{n-1})}_{s_{l}}-\hat{\mathcal{E}}^{(\rho_{n-1})}_{s_{l}}\right\rVert_{\rm Tr}\ ,$ (103)
    where the second inequality follows from comparing Eqs. (77) and (93). Making use of the explicit expressions in Eq. (87) again, the eigenvalues of $(\hat{\mathbf{E}}^{(\rho)}_{s}-\hat{\mathcal{E}}^{(\rho)}_{s})^{2}$ are evaulated to be degenerate and of order $\lambda=\mathcal{O}(s^{4})$. Numerical maximization over all possible values of $s_{l}$ and $\rho_{n-1}$ gives $\frac{1}{2}\lVert\hat{\mathbf{E}}^{(\rho_{n-1})}_{s_{l}}-\hat{\mathcal{E}}^{(\rho_{n-1})}_{s_{l}}\rVert_{\rm Tr}<0.71s_{l}^{2}$. Each $s_{l}$ can be bounded as $\lvert s_{l}\rvert=\lvert\frac{\alpha_{l}L}{M}\rvert\leq\frac{\pi L}{M}$, which leads to a bound
    $\frac{1}{2}\left\lVert\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})\right\rVert_{1}<\frac{7L^{2}}{M}\ .$ (104)
@@
    $\frac{1}{2}\left\lVert\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}(\rho_{n-1})\right\rVert_{1}\leq\frac{1}{2}\lVert\psi_{n-1}-\rho_{n-1}\rVert_{1}=\frac{x_{n-1}}{2}\leq\frac{\epsilon}{2L\pi}\ .$ (105)
@@
    $\frac{1}{2}\lVert\phi_{n}-\rho^{\prime}_{n}\rVert_{1}<\frac{7L^{2}}{M}+\frac{\epsilon}{2L\pi}$ (106)
Combining Eqs. (101) and (105), we achieve
$\frac{1}{2}\lVert\tau-\rho^{\prime}_{n}\rVert_{1}<\delta_{n}-\frac{\epsilon}{2}+\frac{7L^{2}}{M}+\frac{\epsilon}{2L\pi}.$ (107)
@@
$\frac{7L^{2}}{M}+\frac{\epsilon}{2L\pi}<\frac{\epsilon}{4}\ ,$ (108)
which proves Eq. (92). The number of $\rho_{n-1}$ needed for the implementation is $M=\mathcal{O}(L^{2}\epsilon^{-1})$, which concludes the proof. ∎

### v2

@@ IV.2.1 Distance bounds after a DME implementation
Let $\rho=(1-2x)\psi+x\mathmybb{1}_{\rm rel}$ be a state with an associated pure state $\psi$ and a mixedness parameter $x\in[0,\varepsilon]$ for some $0\leq\varepsilon<\frac{1}{2}$. A dynamic unitary channel
$\hat{\mathbf{\Gamma}}^{[\rho]}_{L}=\hat{\mathbf{E}}^{[\rho]}_{\alpha_{L}}\circ\hat{\mathbf{E}}^{[\tau]}_{\beta_{L}}\circ\cdots\circ\hat{\mathbf{E}}^{[\rho]}_{\alpha_{1}}\circ\hat{\mathbf{E}}^{[\tau]}_{\beta_{1}}$ (S93)
that uses the same angles $\{\alpha_{l},\beta_{l}\}_{l}$ used in $\hat{\mathbf{\Gamma}}^{[\psi]}_{L}$ has the distance
$\frac{1}{2}\left\lVert\hat{\mathbf{\Gamma}}^{[\psi]}_{L}-\hat{\mathbf{\Gamma}}^{[\rho]}_{L}\right\rVert_{\rm Tr}\leq L\pi\varepsilon\ ,$ (S94)
@@
$\left\lVert\hat{\mathbf{\Gamma}}^{[\psi]}_{L}-\hat{\mathbf{\Gamma}}^{[\rho]}_{L}\right\rVert_{\rm Tr}\leq\sum_{i=1}^{L}\left\lVert\hat{\mathbf{E}}^{[\psi]}_{\alpha_{i}}-\hat{\mathbf{E}}^{[\rho]}_{\alpha_{i}}\right\rVert_{\rm Tr}\ ,$ (S95)
from the triangle inequality and the unitary invariance of the trace norm. The distance between $\hat{\mathbf{E}}^{[\rho]}_{\alpha}$ and $\hat{\mathbf{E}}^{[\psi]}_{\alpha}$ can be written as
$\frac{1}{2}\left\lVert\hat{\mathbf{E}}^{[\psi]}_{\alpha}-\hat{\mathbf{E}}^{[\rho]}_{\alpha}\right\rVert_{\rm Tr}=\frac{1}{2}\max_{\sigma\in\mathcal{S}(\mathcal{H}_{\rm rel})}\left\lVert\hat{\mathbf{E}}^{[\psi]}_{\alpha}\left(\sigma\right)-\hat{\mathbf{E}}^{[\rho]}_{\alpha}(\sigma)\right\rVert_{1}=\frac{1}{2}\max_{\sigma\in\mathcal{S}(\mathcal{H}_{\rm rel})}\operatorname{Tr}\left\{\sqrt{\left(\hat{\mathbf{E}}^{[\psi]}_{\alpha}\left(\sigma\right)-\hat{\mathbf{E}}^{[\rho]}_{\alpha}(\sigma)\right)^{2}}\right\}\ .$ (S96)
@@
$\psi=\begin{bmatrix}1&0\\ 0&0\end{bmatrix},\quad\rho=\begin{bmatrix}1-x&0\\ 0&x\end{bmatrix},\quad\sigma=\begin{bmatrix}p&q\\ q^{*}&1-p\end{bmatrix}\ ,$ (S97)
@@
$\left(\hat{\mathbf{E}}^{[\psi]}_{\alpha}(\sigma)-\hat{\mathbf{E}}^{[\rho]}_{\alpha}(\sigma)\right)^{2}$ $=\begin{bmatrix}4\lvert q\rvert^{2}\sin^{2}\left(\alpha x\right)&0\\ 0&4\lvert q\rvert^{2}\sin^{2}\left(\alpha x\right)\end{bmatrix}\ .$ (S98)
The state $\sigma$ giving the maximum distance is the one having the maximal $\lvert q\rvert$. The condition that $\sigma\geq 0$ imposes $\lvert q\rvert\leq\frac{1}{2}$, and $\frac{1}{2}\lVert\hat{\mathbf{E}}^{[\psi]}_{\alpha}-\hat{\mathbf{E}}^{[\rho]}_{\alpha}\rVert_{\rm Tr}=\lvert\sin\left(\alpha x\right)\rvert$. From $\alpha\in[-\pi,\pi]$ and $\lvert\alpha\varepsilon\rvert<\frac{\pi}{2}$, we bound $\lvert\sin(\alpha x)\rvert<\pi\varepsilon$ and achieve Eq. (S94). ∎
@@
$\delta_{n}\coloneq\left(T_{(2L+1)}\left(\delta_{n-1}^{-1}\right)\right)^{-1}+\epsilon\ ,$ (S99)
and use it as the parameter $q=\delta_{n}$ when choosing angles Eq. (S74). At the same time, we also demonstrate that $\delta_{n}$ bounds the distance to the target state, after an iteration of DME implementation with error $\epsilon$, i.e. it is the counterpart to Eq. (S76) used for the exact iteration. Precise statements of this result are delineated in Lemma 17 and Corollary 18. Further analysis on Eq. (S99) itself will be made in Sec. IV.3.
@@
We consider iteration $n$ of the nested Grover search algorithm. Assume that the initial state for the iteration $\rho_{n-1}=(1-2x_{n-1})\psi_{n-1}+x_{n-1}\mathmybb{1}_{\rm rel}$ is unknown to us but satisfies the following:
@@
  • The mixedness parameter is bounded by $x_{n-1}\leq\frac{\epsilon}{2L\pi}$ for some number $\epsilon$.
Then, it is possible to prepare a state $\rho^{\prime}_{n}=(1-2x_{n}^{\prime})\psi_{n}+x^{\prime}\mathmybb{1}_{\rm rel}$, such that
$\frac{1}{2}\lVert\tau-\rho^{\prime}_{n}\rVert_{1}$ $\leq\delta_{n}-\frac{\epsilon}{4},$ (S100)
where $\delta_{n}$ is defined by the recurrence relation Eq. (S99). The preparation requires consuming $\mathcal{O}(L^{2}\epsilon^{-1})$ copies of $\rho_{n-1}$.
@@
The preparation is performed by implementing nested fixed-point Grover search with DME memory-usage queries. Firstly, we choose the angles $\{\alpha_{l}(q),\beta_{l}(q)\}_{l}$ for the dynamic unitary $\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}$. Following Lemma 13, we set $q=\operatorname{sech}((2L+1)\operatorname{arcsech}(\delta_{n-1}))$, or equivalently, $q=\delta_{n}-\epsilon$. Secondly, we denote the total number of DME queries ^s$\hat{\mathcal{E}}^{[\rho_{n-1}]}_{s}$ used for implementing $\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}$ as $M$. Since $L$ memory-calls (partial reflections around $\rho_{n-1}$) are contained in $\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}$, each memory-call employs $\frac{M}{L}$ DME memory-usage queries. The implemented channel reads
$\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}\coloneq\left(\hat{\mathcal{E}}^{[\rho_{n-1}]}_{s_{L}}\right)^{\frac{M}{L}}\circ\hat{\mathbf{E}}^{[\tau]}_{\beta_{L}}\circ\cdots\circ\left(\hat{\mathcal{E}}^{[\rho_{n-1}]}_{s_{1}}\right)^{\frac{M}{L}}\circ\hat{\mathbf{E}}^{[\tau]}_{\beta_{1}}\ ,$ (S101)
@@
$\rho^{\prime}_{n}:=\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}\left(\rho_{n-1}\right)=(1-2x_{n}^{\prime})\psi_{n}+x_{n}^{\prime}\mathmybb{1}_{\rm rel}\ .$ (S102)
We additionally define the state after an exact unitary channel $\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}$ as
$\tilde{\rho}_{n}:=\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}\left(\rho_{n-1}\right)=(1-2x_{n-1})\phi_{n}+x_{n-1}\mathmybb{1}_{\rm rel}\ .$ (S103)
Since $\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}$ is a unitary channel, it leaves the projector $\mathmybb{1}_{\rm rel}$ intact in Eq. (S103), and the associated pure state
$\phi_{n}=\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})\ ,$ (S104)
@@
    $\lVert\tau-\phi_{n}\rVert_{1}=\left\lVert\tau-\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}\leq\left\lVert\tau-\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}+\left\lVert\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}(\psi_{n-1})-\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}\ .$ (S105)
    The second term in the RHS is bounded by the trace norm distance of channels $\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}$ and $\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}$. Applying Proposition 16 and using the assumption $x_{n-1}\leq\frac{\epsilon}{2L\pi}$,
    $\left\lVert\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}(\psi_{n-1})-\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}\leq\left\lVert\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}-\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}\right\rVert_{\rm Tr}\leq\epsilon\ .$ (S106)
    The other term $\lVert\tau-\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}(\psi_{n-1})\rVert_{1}$ resembles the final distance after an exact nested fixed-point Grover search in Lemma 13. However, we cannot directly apply Lemma 13, since we do not know the initial distance $\lVert\tau-\psi_{n-1}\rVert_{1}$ exactly; we only assume that it is bounded by $\delta_{n-1}$. Hence, we must start from the more general relation Eq. (S75), which gives the equality
    $\frac{1}{2}\left\lVert\tau-\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}=(\delta_{n}-\epsilon)T_{2L+1}\left(T_{1/(2L+1)}\left((\delta_{n}-\epsilon)^{-1}\right)\frac{1}{2}\left\lVert\tau-\psi_{n-1}\right\rVert_{1}\right)\ ,$ (S107)
    substituting $q=\delta_{n}-\epsilon$. Recalling that $\delta_{n}-\epsilon=(T_{(2L+1)}(\delta_{n-1}^{-1}))^{-1}$ and Eq. (S71), we get $T_{1/(2L+1)}((\delta_{n}-\epsilon)^{-1})=\delta_{n-1}^{-1}$. Eq. (S107) then becomes
    $\frac{1}{2}\left\lVert\tau-\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}=(\delta_{n}-\epsilon)T_{2L+1}\left(\delta_{n-1}^{-1}\frac{1}{2}\left\lVert\tau-\psi_{n-1}\right\rVert_{1}\right)\ .$ (S108)
    Chebyshev polynomials follow $T_{2L+1}(x)\leq 1$ whenever $x\in[0,1]$, and the argument of $T_{2L+1}$ in the RHS, $\delta_{n-1}^{-1}\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}$ is indeed smaller than or equal to $1$ from the assumption of the Lemma. Then $\frac{1}{2}\lVert\tau-\hat{\mathbf{\Gamma}}^{[\psi_{n-1}]}_{L}(\psi_{n-1})\rVert_{1}\leq\delta_{n}-\epsilon$ follows, and we finally obtain the bound
    $\frac{1}{2}\lVert\tau-\phi_{n}\rVert_{1}\leq\delta_{n}-\frac{\epsilon}{2}\ .$ (S109)
@@
    $\lVert\phi_{n}-\rho^{\prime}_{n}\rVert_{1}$ $\leq\left\lVert\phi_{n}-\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}+\left\lVert\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})-\rho^{\prime}_{n}\right\rVert_{1}$
    $=\left\lVert\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}+\left\lVert\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\rho_{n-1})\right\rVert_{1}\ .$ (S110)
@@
    $\left\lVert\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}\leq\left\lVert\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}-\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}\right\rVert_{\rm Tr}\leq\sum_{l=1}^{L}\frac{M}{L}\left\lVert\hat{\mathbf{E}}^{[\rho_{n-1}]}_{s_{l}}-\hat{\mathcal{E}}^{[\rho_{n-1}]}_{s_{l}}\right\rVert_{\rm Tr}\ ,$ (S111)
    where the second inequality follows from comparing Eqs. (S93) and (S101). Making use of the explicit expressions in Eq. (S97) again, the eigenvalues of $(\hat{\mathbf{E}}^{[\rho]}_{s}-\hat{\mathcal{E}}^{[\rho]}_{s})^{2}$ are evaluated to be degenerate and of order $\lambda=\mathcal{O}(s^{4})$. Numerical maximization over all possible values of $s_{l}$ and $\rho_{n-1}$ gives $\frac{1}{2}\lVert\hat{\mathbf{E}}^{[\rho_{n-1}]}_{s_{l}}-\hat{\mathcal{E}}^{[\rho_{n-1}]}_{s_{l}}\rVert_{\rm Tr}<0.71s_{l}^{2}$. Each $s_{l}$ can be bounded as $\lvert s_{l}\rvert=\lvert\frac{\alpha_{l}L}{M}\rvert\leq\frac{\pi L}{M}$, which leads to a bound
    $\frac{1}{2}\left\lVert\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})\right\rVert_{1}\leq\frac{1}{2}\left\lVert\hat{\mathbf{\Gamma}}^{[\rho_{n-1}]}_{L}-\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}\right\rVert_{\rm Tr}<\frac{7L^{2}}{M}\ .$ (S112)
@@
    $\frac{1}{2}\left\lVert\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\psi_{n-1})-\hat{\mathmybb{\Gamma}}^{[\rho_{n-1}]}_{L}(\rho_{n-1})\right\rVert_{1}\leq\frac{1}{2}\lVert\psi_{n-1}-\rho_{n-1}\rVert_{1}=x_{n-1}\leq\frac{\epsilon}{2L\pi}\ .$ (S113)
@@
    $\frac{1}{2}\lVert\phi_{n}-\rho^{\prime}_{n}\rVert_{1}<\frac{7L^{2}}{M}+\frac{\epsilon}{2L\pi}$ (S114)
Combining Eqs. (S109) and (S113), we achieve
$\frac{1}{2}\lVert\tau-\rho^{\prime}_{n}\rVert_{1}<\delta_{n}-\frac{\epsilon}{2}+\frac{7L^{2}}{M}+\frac{\epsilon}{2L\pi}.$ (S115)
@@
$\frac{7L^{2}}{M}+\frac{\epsilon}{2L\pi}<\frac{\epsilon}{4}\ ,$ (S116)
which proves Eq. (S100). The number of $\rho_{n-1}$ needed for the implementation is $M=\mathcal{O}(L^{2}\epsilon^{-1})$, which concludes the proof. ∎